# VulMCI : Code Splicing-based Pixel-row Oversampling for More Continuous Vulnerability Image Generation

### v1

@@ Abstract.
In recent years, the rapid development of deep learning technology has brought new prospects to the field of vulnerability detection. Many vulnerability detection methods involve converting source code into images for detection, yet they often overlook the quality of the generated images. Due to the fact that vulnerability images lack clear and continuous contours, unlike images used in object detection, Convolutional Neural Networks (CNNs) tend to lose semantic information during the convolution and pooling processes. Therefore, this paper proposes a pixel row oversampling method based on code line concatenation to generate more continuous code features, addressing the issue of discontinuity in code image coloration.Building upon these contributions, we propose the vulnerability detection system VulMCI and conduct tests on the SARD and NVD datasets. Experimental results demonstrate that VulMCI outperforms seven state-of-the-art vulnerability detectors (namely Checkmarx, FlawFinder, RATS, VulDeePecker, SySeVR, VulCNN, and Devign). Compared to other image-based methods, VulMCI shows improvements in various metrics, including a 2.877% increase in True Positive Rate (TPR), a 5.446% increase in True Negative Rate (TNR), and a 5.91% increase in Accuracy (ACC). On the NVD real-world dataset, VulMCI achieves an average accuracy of 5.162%, confirming its value in practical vulnerability detection applications. We will share the source code of the paper on GitHub after it is accepted.
1. Introduction
@@
With the widespread use of computer networks and the ubiquity of the Internet, vulnerability detection has become increasingly crucial as malicious hackers and cybercriminals continuously seek novel methods to infiltrate systems and pilfer sensitive information. Furthermore, as software scales and complexities continue to grow, vulnerabilities within source code have become more insidious and pervasive, posing a severe threat to information security. According to the latest data from the National Vulnerability Database (NVD)(of Standards and Technology, 2023), we can observe that in 2021, the number of disclosed vulnerabilities surpassed the twenty-thousand mark, and even though the figure decreased to slightly over thirteen thousand in 2022, it remains a substantial quantity. Consequently, the need for efficient and automated source code vulnerability detection has emerged as an urgent requirement to ensure the robustness and reliability of software systems.
@@
Traditional vulnerability detection methods typically employ techniques based on code similarity(Jang et al., 2012; Kim et al., 2017; Li and Ernst, 2012; Li et al., 2016; Pham et al., 2010) or rule-based approaches(FlawFinder, 2021; Rough Audit Tool for Security, 2021; Checkmarx, 2021) to analyze source code in order to identify potential vulnerabilities. However, these conventional methods exhibit several notable limitations. Firstly, they often require a significant amount of manual work and rule formulation, which becomes impractical when dealing with large-scale code repositories. Secondly, they frequently struggle to address complex vulnerability types, especially those with concealment and variants. Finally, these methods tend to produce false positives and false negatives when handling extensive code repositories, thus compromising the accuracy and efficiency of the detection process.
@@
Existing deep learning-based vulnerability detection systems utilize various methods, including processing code into textual representations using natural language processing(Li et al., 2018; Russell et al., 2018b), analyzing code structure graphs using graph neural networks(Duan et al., 2019a; Zhou et al., 2019), and generating images from code for detection(Wu et al., 2022). While each of these methods has its merits, there is still room for improvement. In some prior studies(Li et al., 2018, 2022), custom Common Weakness Enumeration (CWE) vulnerability sets were used to match and locate vulnerabilities. However, this approach may lead to incomplete vulnerability coverage and instances where vulnerability labels do not match in real-world datasets. Additionally, many methods(Li et al., 2022; Lv et al., 2023) rely on the concept of code slicing to eliminate redundant information, depending not only on highly accurate vulnerability line localization but also risking the loss of complete program semantic information during slicing.
Furthermore, methods that transform source code into images, such as VulCNN(Wu et al., 2022), although preserving the complete semantics and structural information of functions, have been found through experiments to generate RGB images with significant instances of black and discontinuous scattered color points in the channels. This characteristic is considered detrimental to the classification task of Convolutional Neural Network (CNN) models.
In this paper, we adopt the baseline method of converting source code into images and make improvements to address the issue of discontinuity in image coloration. It is important to note that we do not simply apply morphological operations to the images, as doing so would disturb the weight differences between sentences. Instead, we apply a pixel row oversampling algorithm to source code lines with dependencies to enhance the continuity between them. The main advantage of this approach is the reduction of data noise, thereby minimizing differences between adjacent data points. This results in a more stable model that is less sensitive to small variations in the input data, contributing to improved model stability and generalization capabilities.
@@
We identified a novel issue wherein there are discontinuous and abrupt numerical changes when code is transformed into image representations. This discontinuity can potentially have a negative impact on the training of Convolutional Neural Network (CNN) models, particularly when a significant amount of line information is lost after the image undergoes pooling operations.
We propose a pixel row oversampling algorithm based on code concatenation. To address the issue of discontinuous color points in images directly generated from code, we employ an adaptive segmentation method to insert concatenated rows with weights between adjacent rows. This approach balances local variations in the image, highlighting important features and enhancing the CNN’s ability to extract key features, ultimately contributing to improved classification performance.
We introduce a new system, VulMCI, which can extract function samples from complex source code and generate code feature images with more continuous upper and lower pixel rows.
We introduce a new hypothesis that high-quality vulnerability detection images can enhance the detection performance of Convolutional Neural Network (CNN).
Paper organization. The remaining sections of this paper are organized as follows. Section 2 introduces our method. Section 3 presents the experimental results. Section 4 discusses future work. Section 5 provides an overview of related work. Finally, Section 6 summarizes the key findings of this paper.
2. Our Method
The existing methods for generating images from code overlook the continuity of pixel rows, resulting in suboptimal image quality. Therefore, we propose a systematic approach in VulMCI to generate high-quality images with more continuous upper and lower pixel rows. As illustrated in Figure1, VulMCI comprises four main stages: building function samples, pixel-row oversampling, code encoding and embedding, and generating three-channel images.
@@ 2. Our Method
Pixel-row oversampling: In existing functions, we employ an adaptive segmentation method to insert new concatenated rows between adjacent rows.
Generating three-channel images: We adopt three centrality analysis indicators from VulCNN(Wu et al., 2022) (degree centrality(Freeman, 1978), katz centrality(Katz, [n. d.]), closeness centrality(Freeman, 1978)). Different from VulCNN, which generates Program Dependency Graph (PDG), we generate Code Property Graph (CPG) for functions, additionally incorporating Abstract Syntax Tree (AST) and Control Dependency Graph (CDG).
2.1. Pixel-row oversampling
Given adjacent code lines, denoted as Line and Line (where represents the line number), adaptive concatenation is performed between these two adjacent lines. The concatenation process is carried out as follows:
Step 1: In Line , take the number obtained from Formula 1 as the length of the substring, and obtain the substring .
(1) -
Step 2: In Line , take the number obtained from Formula 2 as the length of the substring, and obtain the substring .
(2) -
Step 3: Obtain the new concatenated line , and complete the words at the character-level separation points.
For , k-1 new concatenated lines will be gradually inserted between adjacent lines. For example, when , the values for are 1 and 2, and two concatenated lines will be inserted between each adjacent pair of lines.
In the red-boxed portion of the function samples shown in Figure 1, where , (including the number of spaces between words), , , and , , thus only one concatenated line will be inserted. For ,= and &=.
, . The resulting concatenated line is .
2.2. Generating a three-channel image
We inputted function samples into the Joern tool(Yamaguchi et al., 2014), which utilizes static analysis to generate a Code Property Graph (CPG). The CPG includes Control Flow Graph (CFG), Control Dependency Graph (CDG), Data Dependency Graph (DDG), and Abstract Syntax Tree (AST). Similar to VulCNN, we derived an adjacency matrix and calculated in-degrees, out-degrees, and degrees for all code lines based on the edge and node information in CPG. Subsequently, we conducted centrality analysis on the CPG, computing the degree centrality, Katz centrality, and closeness centrality for each code line node. The formulas for these three centrality metrics are as follows, following similar procedures as in VulCNN.
Degree Centrality(Freeman, 1978):
(3) Where is the degree of node , and is the number of nodes in the graph.
Katz Centrality(Katz, [n. d.]):
(4) Where and are constants, is the adjacency matrix of the graph, and is the centrality of the -th node.
(5) Where is the largest eigenvalue of the adjacency matrix .
Closeness Centrality(Freeman, 1978):
(6) Where is the distance between nodes and , and is the number of nodes in the graph.
The formula for the centrality metrics of the newly generated concatenated line is as follows:
2.3. algorithmic descriptions
Algorithm 1 outlines how VulMCI transforms function code into a three-channel image. Firstly, we preprocess the function code, and then we perform adaptive segmentation based on to concatenate and insert new lines between adjacent code lines, forming new lines. Subsequently, we calculate centrality metrics for each code line based on the Code Property Graph (CPG). Following this, each of these code lines is vectorized using sent2vec, producing corresponding vector representations. These vectors are then multiplied by three expanded lists of centrality metrics and mapped into three pixel channels resembling RGB. Ultimately, this process creates a code image with improved continuity.
2.4. Theoretical Analysis
Definition 2.0 ().
@@ 2.4. Theoretical Analysis
Step 1: We demonstrate the continuity between adjacent pixel rows and concatenated pixel rows under specific conditions. The objective of the sent2vec(Pagliardini et al., 2018) model is to learn word and sentence embedding by minimizing a loss function. The simplified representation of the objective is as follows:
@@
The complete training objective of the Sent2Vec(Pagliardini et al., 2018) model, an unsupervised learning method for learning and inferring sentence embeddings, is as follows:
@@
The first part of the loss function focuses on positive samples, encouraging similar sentences to be closer in the embedding space through the logistic regression loss function . The second part addresses negative samples, pushing dissimilar sentences further apart in the embedding space through the logistic regression loss function . The overall objective is to minimize the loss function by learning and , resulting in embeddings with enhanced semantic relationships between sentences and words. By concatenating adjacent code tokens and inserting them in between, more context information is introduced, making adjacent code lines closer in the embedding space. The encoding and embedding stage in Figure 1 illustrates this characteristic. Thus, the training objective of sent2vec yields a set of UV parameters, theoretically resulting in a small gap. This demonstrates the continuity between adjacent vector rows in our approach. ∎
2.5. Classification
@@
The code length distribution of function samples in the experimental dataset (SARD) is illustrated in Figure 2. It is observed that the majority of function lengths are less than 100 lines. Therefore, we have chosen a code length threshold of 100 lines to generate our input images. And previous experiments have shown that choosing 100 lines as a threshold can balance overhead and detection performance well. Furthermore, we dynamically adjust the parameter for the concatenated lines, setting the code line threshold to .We take 100 lines as an example. For functions with code lengths less than 100, we pad the vectors with zeros at the end. In the case of functions exceeding 100 lines, we truncate the trailing portion of the vectors. The input image size is set to 3×100×128, where 3 corresponds to three channels (i.e., ”degree channel,” ”Katz channel,” and ”proximity channel”), 100 is the code line threshold, and 128 represents the dimensionality of the sentence vectors.
After generating images of fixed size, we build a Convolutional Neural Network (CNN) model through training on these images. As depicted in Figure 3, we employ convolutional filters of varying shapes, with dimensions m * 128, to ensure each filter can extract features across the entire space of the embedded sentences. The size of the filters roughly determines the length of the sentence sequences considered simultaneously. In VulMCI, we select 10 filters of different sizes (from 1 to 10), each containing 32 feature maps to capture features from different parts of the image. Following the max-pooling operation, the length of our fully connected layer is 320. Table 1 provides a detailed description of the parameters used in VulMCI. The entire model adopts Rectified Linear Unit (ReLU)(Dahl et al., 2013) as the non-linear activation function. Additionally, we utilize cross-entropy loss as the loss function in CNN for penalizing incorrect classifications. We employ Adam(Kingma and Ba, 2014) as the optimizer with a learning rate set to 0.001. Once the training is completed, we use the trained CNN model to classify new functions, determining whether they possess vulnerabilities.
@@ 3. EXPERIMENTS
RQ1: How much does the detection performance improve with different values of the parameter in the pixel-row oversampling operation?
@@
3.1. Experiment Settings
The dataset used in this paper is sourced from the vulnerability dataset provided by SySeVR(Li et al., 2022). This dataset originates from SARD(sar, 2018) and NVD(of Standards and Technology, 2023), encompassing a total of 126 vulnerability types, each uniquely identified by a Common Weakness Enumeration (CWE) ID(cwe, 2018). The SARD dataset comprises production, synthetic, and academic programs (referred to as test cases) categorized as ”good” (i.e., without vulnerabilities), ”bad” (i.e., containing vulnerabilities), and ”mixed” (i.e., vulnerabilities with available patch versions). The NVD dataset includes vulnerabilities in 19 popular C/C++ open-source products (software systems), along with possible diff files describing the differences between susceptible code and its patched versions. The SARD dataset contains a total of 21,233 program files, while the NVD dataset includes 2,011 real vulnerabilities and their corresponding fixed version files. Our system processed the dataset by extracting function samples from program files and removing function samples with less than ten lines of code. This exclusion was based on the observation that the majority of functions with fewer than ten lines only involve external function calls and cannot be easily classified as usable function samples. Consequently, our final processing resulted in 13,165 samples of vulnerable functions and 5,612 samples of non-vulnerable functions.
In this study, we employed a k-fold cross-validation (k=5) strategy to partition the dataset, creating five mutually exclusive subsets. Four subsets were used for training the model, while one subset was reserved for testing. This partitioning approach ensures the utilization of distinct data subsets for both training and testing phases, facilitating the assessment of the model’s generalization performance. The devices and experimental parameters employed in this study are presented in Tables 1 and 2, respectively. To comprehensively evaluate the performance of our proposed method, we utilized multiple evaluation metrics, including False Positive Rate (FPR), False Negative Rate (FNR), Precision (Pr), Recall (Re), F1 Score (F1), and Accuracy (ACC). Each metric provides detailed insights into the algorithm’s performance across different aspects, offering a nuanced understanding of its strengths and limitations.

### v2

@@ Abstract
In recent years, the rapid development of deep learning technology has brought new prospects to the field of vulnerability detection. Many vulnerability detection methods transform source code into images for detection, but they neglect the issue of image quality generation. Since vulnerability images do not possess clear and continuous contours like object detection images, Convolutional Neural Networks (CNNs) tend to lose semantic information during the convolution and pooling process. Therefore, this paper proposes a pixel row oversampling method based on code line concatenation to generate more continuous code features, addressing the problem of discontinuous colors in code images. The feasibility of the proposed method is theoretically analyzed and verified. Based on these efforts, we introduce the vulnerability detection system VulMCI, which is tested on two datasets, SARD and NVD. Experimental results demonstrate that VulMCI outperforms eight state-of-the-art vulnerability detectors (i.e., Checkmarx, FlawFinder, RATS, VulDeePecker, SySeVR, Devign, VulCNN, and AMPLE) in accuracy on the SARD dataset. Compared to other image-based methods, VulMCI shows improvements in all metrics, including a 19.82% increase in True Positive Rate (TPR), a 4.95% increase in True Negative Rate (TNR), and an 8.89% increase in accuracy (ACC). On the NVD real-world dataset, VulMCI achieves an average accuracy of 90.41%.
Index Terms:
Vulnerability Detection, Scurity, Deep Learning, Program Analysis, Program RepresentationI Introduction
@@
With the widespread use of computer networks and the ubiquity of the Internet, vulnerability detection has become increasingly crucial as malicious hackers and cybercriminals continuously seek novel methods to infiltrate systems and pilfer sensitive information. Furthermore, as software scales and complexities continue to grow, vulnerabilities within source code have become more insidious and pervasive, posing a severe threat to information security. According to the latest data from the National Vulnerability Database (NVD)[1], we can observe that in 2021, the number of disclosed vulnerabilities surpassed the twenty-thousand mark, and even though the figure decreased to slightly over thirteen thousand in 2022, it remains a substantial quantity. Consequently, the need for efficient and automated source code vulnerability detection has emerged as an urgent requirement to ensure the robustness and reliability of software systems.
@@
Traditional vulnerability detection methods typically employ techniques based on code similarity[2, 3, 4, 5, 6] or rule-based approaches[7, 8, 9] to analyze source code in order to identify potential vulnerabilities. However, these conventional methods exhibit several notable limitations. Firstly, they often require a significant amount of manual work and rule formulation, which becomes impractical when dealing with large-scale code repositories. Secondly, they frequently struggle to address complex vulnerability types, especially those with concealment and variants. Finally, these methods tend to produce false positives and false negatives when handling extensive code repositories, thus compromising the accuracy and efficiency of the detection process.
@@
Existing deep learning-based vulnerability detection systems utilize various methods, including processing code into textual representations using natural language processing[10, 11], analyzing code structure graphs using graph neural networks[12, 13], and generating images from code for detection[14]. While each of these methods has its merits, there is still room for improvement. In some prior studies[10, 15], custom Common Weakness Enumeration (CWE) vulnerability sets were used to match and locate vulnerabilities. However, this approach may lead to incomplete vulnerability coverage and instances where vulnerability labels do not match in real-world datasets. Additionally, many methods[15, 16] rely on the concept of code slicing to eliminate redundant information, depending not only on highly accurate vulnerability line localization but also risking the loss of complete program semantic information during slicing.
Furthermore, methods that transform source code into images, such as VulCNN[14], although preserving the complete semantics and structural information of functions, have been found through experiments to generate RGB images with significant instances of black and discontinuous scattered color points in the channels. This characteristic is considered detrimental to the classification task of Convolutional Neural Network (CNN) models.
In this paper, we adopt the baseline method of converting source code into images and propose improvements to address the issue of discontinuity between rows of pixels in the images. We apply a pixel row oversampling algorithm to the source code lines with dependencies to enhance the continuity between code lines. The main advantage of this approach is to reduce data noise and minimize differences between adjacent data points, thereby making the model more robust to small variations in the input data. This contributes to improving model stability and generalization capability.
@@
We present a novel framework for vulnerability image detection called VulMCI, which incorporates a method of pixel row oversampling using code control flow. This method generates code feature images with more continuous rows of pixels, effectively enhancing the identification of vulnerability features.
We propose a pixel row oversampling algorithm based on code concatenation. To address the issue of discontinuous numerical values in directly generated code images, we utilize the relationships between nodes in the control flow graph to insert new code lines between adjacent rows. This balances local variations in the image and highlights important features, thereby improving the CNN’s ability to extract key features and enhancing classification performance. s
We present a novel finding that when code is converted into image representation, there exist discontinuous and abrupt numerical changes. This discontinuity may negatively impact the training of Convolutional Neural Network (CNN) models, especially after the images undergo pooling operations, leading to the loss of significant row information.
Paper organization. The remaining sections of this paper are organized as follows. Section 2 presents the motivation of our paper. Section 3 introduces our method. Section 4 presents the experimental results. Section 5 discusses future work. Section 6 provides an overview of related work. Finally, Section 7 summarizes the key findings of this paper.
II Motivation
Figure 1 depicts a vulnerability code image generated using our VulCNN[14] method. We observed the presence of numerous redundant zero vectors and discontinuous color points, which are detrimental to the subsequent training of CNN models. This is attributed to the significant differences in feature representation between vulnerability detection tasks and object detection tasks when using images for classification. Object detection tasks typically require models to capture continuous edge features of target objects, which are retained to some extent after multiple convolution and pooling operations, as illustrated in Figure 2. However, in vulnerability detection tasks, the features of key statements often occupy a relatively small proportion of the data, and their distribution may be more dispersed. This leads to significant fluctuations in pixel values and the generation of discontinuous color points when directly converting vulnerability code into images. This scenario presents two main challenges for classical convolutional neural networks (CNNs) when processing vulnerability code images:
1.
Information Loss: Vulnerability code images are likely to lose crucial semantic information after convolution and pooling operations, especially concerning the features of key vulnerability statements. This poses difficulties in training and inference for vulnerability detection models.
2.
Interference Noise: In addition to key statements, vulnerability code images typically contain a large number of irrelevant statements, which may introduce additional interference noise and reduce the model’s accurate understanding of vulnerability features.
In response to these challenges, we propose a pixel-row oversampling method based on control flow graph-guided code splicing. By integrating the structured features of the code, we selectively splice correlated code lines to generate new sample rows. This approach aims to bridge code segments with flow relationships, enhancing contextual continuity and thereby improving the extraction of vulnerability features.
III Our Method
Existing methods for generating images from code overlook the continuity of pixel rows, resulting in low-quality images. Therefore, we propose the system approach of VulMCI to generate high-quality images with more continuous rows. As illustrated in Figure 3, VulMCI consists of four main stages: constructing function samples, pixel row oversampling, code encoding and embedding, and generating grayscale images.
@@ III Our Method
Pixel-row oversampling: We generate Code Property Graphs (CPGs) from the extracted function samples, and utilize the edge-node relationships of Control Flow Graphs (CFGs) to synthesize new code lines inserted into the source code, where each node represents a line of code.
Generating three-channel images: We arrange the generated code line vectors together to form a grayscale image of vulnerabilities, which is used for subsequent training and classification by CNNs.
III-A Pixel-row oversampling
We utilize Joern[18] to generate Code Property Graphs (CPGs) from the extracted function samples, and then splice them based on the node relationships of Control Flow Graphs (CFGs). The control flow graph describes the control flow of the program, representing the transitions of control flow during program execution. Thus, splicing based on CFG node relationships enhances contextual connections. The splicing process follows these steps: we first traverse the line numbers of the code, and if a node corresponding to the line number has edges of CFG type, we concatenate the code of that node with the code of the target node completely, inserting it as a new code sample line after the node line. If the node corresponding to the code line does not belong to CFG, no splicing is performed, but the code line content is retained. As shown in the red-boxed portion of the function samples in Figure 3, lines 4 and 5 correspond to nodes with existing edges in the CFG. Therefore, the codes of lines 4 and 5 are concatenated and inserted between the two lines. Finally, the function samples after pixel row oversampling are vectorized using sent2vec[17], where each line of code is embedded as a row vector. These vector rows are then arranged together to generate a grayscale image. Figure3 and Algorithm 1 describe how VulMCI converts function code into a grayscale image.
III-B Theoretical Analysis
Definition 1.
@@ III-B Theoretical Analysis
Step 1: We demonstrate the continuity between adjacent pixel rows and concatenated pixel rows under specific conditions. The objective of the sent2vec[17] model is to learn word and sentence embedding by minimizing a loss function. The simplified representation of the objective is as follows:
@@
The complete training objective of the Sent2Vec[17] model, an unsupervised learning method for learning and inferring sentence embeddings, is as follows:
@@
The first part of the loss function focuses on positive samples, encouraging similar sentences to be closer in the embedding space through the logistic regression loss function . The second part addresses negative samples, pushing dissimilar sentences further apart in the embedding space through the logistic regression loss function . The overall objective is to minimize the loss function by learning and , resulting in embeddings with enhanced semantic relationships between sentences and words. By splicing and inserting adjacent code nodes into the middle based on the control flow graph, we introduce more contextual information, making contextually relevant code lines closer in the embedding space. This characteristic is reflected in the code encoding and embedding stage shown in Figure3. Therefore, the training objective of sent2vec[17] can obtain a set of UV parameters, theoretically leading to a smaller gap, proving the continuity between adjacent vector rows. ∎
III-C Classification
@@
The code length distribution of function samples in the experimental dataset (SARD) is illustrated in Figure 4. It is observed that the majority of function lengths are less than 100 lines. Therefore, we have chosen a code length threshold of 100 lines to generate our input images. And previous experiments have shown that choosing 100 lines as a threshold can balance overhead and detection performance well. For functions with code lengths less than 100, we pad the vectors with zeros at the end. In the case of functions exceeding 100 lines, we truncate the trailing portion of the vectors. The input image size is set to 100×128, 100 is the code line threshold, and 128 represents the dimensionality of the sentence vectors.
After generating images of fixed size, we build a Convolutional Neural Network (CNN) model through training on these images. As depicted in Figure 5, we employ convolutional filters of varying shapes, with dimensions m * 128, to ensure each filter can extract features across the entire space of the embedded sentences. The size of the filters roughly determines the length of the sentence sequences considered simultaneously. In VulMCI, we select 10 filters of different sizes (from 1 to 10), each containing 32 feature maps to capture features from different parts of the image. Following the max-pooling operation, the length of our fully connected layer is 320. Table I provides a detailed description of the parameters used in VulMCI. The entire model adopts Rectified Linear Unit (ReLU)[19] as the non-linear activation function. Additionally, we utilize cross-entropy loss as the loss function in CNN for penalizing incorrect classifications. We employ Adam[20] as the optimizer with a learning rate set to 0.001. Once the training is completed, we use the trained CNN model to classify new functions, determining whether they possess vulnerabilities.
@@ IV EXPERIMENTS
RQ1: What are the advantages of VulMCI’s pixel row oversampling method compared to other oversampling methods?
@@
IV-A Experiment Settings
The dataset used in this paper is sourced from the vulnerability dataset provided by SySeVR[15]. This dataset originates from SARD[21] and NVD[1], encompassing a total of 126 vulnerability types, each uniquely identified by a Common Weakness Enumeration (CWE) ID[22]. The SARD dataset comprises production, synthetic, and academic programs (referred to as test cases) categorized as ”good” (i.e., without vulnerabilities), ”bad” (i.e., containing vulnerabilities), and ”mixed” (i.e., vulnerabilities with available patch versions). The NVD dataset includes vulnerabilities in 19 popular C/C++ open-source products (software systems), along with possible diff files describing the differences between susceptible code and its patched versions. The SARD dataset contains a total of 21,233 program files, while the NVD dataset includes 2,011 real vulnerabilities and their corresponding fixed version files. Our system processed the dataset by extracting function samples from program files and removing function samples with less than ten lines of code. This exclusion was based on the observation that the majority of functions with fewer than ten lines only involve external function calls and cannot be easily classified as usable function samples. Ultimately, we obtained 12,116 vulnerability function samples and 4,660 non-vulnerability function samples from the SARD dataset, and 1,049 vulnerability function samples and 952 non-vulnerability function samples from the NVD dataset.
In this study, we first conducted comparative and ablation experiments using the SARD dataset, and then employed the NVD dataset in RQ3 to evaluate the effectiveness of vulnerability detection on real vulnerabilities. We adopted a k-fold cross-validation (k=5) approach to partition the entire dataset into 5 mutually exclusive subsets, with 4 subsets used for model training and 1 subset for model testing. This partitioning approach ensures the utilization of distinct data subsets for both training and testing phases, facilitating the assessment of the model’s generalization performance. The devices and experimental parameters employed in this study are presented in Tables I and II, respectively. To comprehensively evaluate the performance of our proposed method, we utilized multiple evaluation metrics, including False Positive Rate (FPR), False Negative Rate (FNR), Precision (Pr), Recall (Re), F1 Score (F1), and Accuracy (ACC). Each metric provides detailed insights into the algorithm’s performance across different aspects, offering a nuanced understanding of its strengths and limitations.